Mac and Charlie have different abilities in the production of candles and drinks. Mac can produce...

  1. Mac and Charlie have different abilities in the production of candles and drinks. Mac can produce 50 candles per day and 75 drinks per day. Charlie can produce 40 candles per day and 50 drinks per day.

    Mac’s opportunity cost of producing a candle is _____ and Charlie’s opportunity cost of producing a candle is _____.

    1.5 drinks; 4/5 of a drink

    2/3 of a drink; 4/5 of a drink

    1.5 drinks; 1.25 drinks

    None of the above

QUESTION 9

  1. Mac and Charlie have different abilities in the production of candles and drinks. Mac can produce 50 candles per day and 75 drinks per day. Charlie can produce 40 candles per day and 50 drinks per day.

    Who has the absolute advantage in the production of candles?

    Mac

    Charlie

    Not enough information

QUESTION 10

  1. Mac and Charlie have different abilities in the production of candles and drinks. Mac can produce 50 candles per day and 75 drinks per day. Charlie can produce 40 candles per day and 50 drinks per day.

    Who has the absolute advantage in the production of drinks?

    Mac

    Charlie

    Not enough information

QUESTION 11

  1. Mac and Charlie have different abilities in the production of candles and drinks. Mac can produce 50 candles per day and 75 drinks per day. Charlie can produce 40 candles per day and 50 drinks per day.

    Who has the comparative advantage in the production of candles?

    Mac

    Charlie

    Not enough information
Answer #1

8. Option 3. 1.5 drinks; 1.25 drinks

Explanation: Mac's opportunity cost of 1 candle = 75/50 = 1.5 drinks. Charlie’s opportunity cost of producing a candle is = 50/40 = 1.25 drinks.

9. Option 1. Mac

Explanation: Mac can produce more candles than Charlie in a given day.

10. Option 1. Mac

Explanation: Mac can produce more drinks than Charlie in a given day.

11. Option 2. Charlie

Explanation:

Mac's opportunity cost of 1 candle = 75/50 = 1.5 drinks. Charlie’s opportunity cost of producing a candle is = 50/40 = 1.25 drinks.

The person who has a lower opportunity cost is said to have a comparative advantage in a good.